# Schwarzschild Black Hole Surrounded by Perfect Fluid Dark Matter in the presence of Quintessence Matter Field

### v1

@@ 3 Black Hole’s Thermodynamics
Fig. 3 shows that the Hawking temperature decreases in all cases as the event horizon increases similar to the ordinary case. However, only for $\omega_{q}=-0.35$, the Hawking temperature tends to zero at infinite horizon radius as in the standard case. For the other three cases, the Hawking temperature becomes zero at an upper bound horizon value and the PFDM parameter affects that value. In fact, we observe in all cases, for a fixed event horizon radius within a physically meaningful range, black holes with greater dark matter parameters would have higher temperatures.
@@
As we emphasized before, we see that the $\omega_{q}=-0.35$ scenario differs from the other cases. Fig. 5 shows that the heat capacity function takes only negative values and the alteration of the dark matter parameters does not provide a significant modification. These findings mean that in this scenario the black hole is always thermally unstable [173]. In other cases, we observe that the heat capacity is first negative and then positive. This indicates a first-order phase transition at the horizon where specific heat equals zero. In other words, the vanishing value of the specific heat states that the black hole ceases to exchange radiation with its surroundings. Therefore, at that horizon black hole becomes stable and a remnant mass occurs. To clarify these discussions, we calculate the event horizon radii where specific heat disappears and present their values in Table 3.
@@ 4 Black Hole’s Shadow
in terms of $r_{ph}$. After we substitute the lapse function, Eq. (7), we arrive at a subsequent equation
@@ 5.2 Mashhoon approximation
In [186], Mashhoon proposed to use Pösch-Teller (PT) potential for calculating the quasinormal modes
@@
Here, $r_{0}^{\ast}$ is the position of the maximum of $\mathcal{V}\left(r^{\ast}\right)$ where $\mathcal{V}\left(r_{0}^{\ast}\right)=V_{0}$. The quasinormal modes of the Pösch-Teller (PT) potential can be computed analytically
@@ 6 Conclusions
In this paper, we handled the Schwarzschild black hole surrounded by PFDM in the presence of the quintessence matter field, and we explored its thermodynamics, shadows, and quasinormal modes. To this end, we initially derived the lapse function of the black hole and discussed the lapse function’s characteristics in four different quintessence field scenarios. We found that the quintessence matter field significantly alters the behavior of the lapse function unless the quintessence state parameter is large enough. Then, we examined the mass function. We observed that dark matter sets a lower limit on the event horizon, but also expands its upper bound value. We noted that achieves a maximum mass value at a larger event horizon with a greater dark matter parameter. Next, we examined the black hole’s thermodynamics. At first, We derived the modified Hawking temperature. We found that it tends to zero unless the quintessence state parameter is not large enough. We showed the event horizon entropy is not affected by dark matter and quintessence matter field. Then. we investigated the heat capacity function. We showed that except very large quintessence state parameter case, the black hole has a first-order phase transition. In these cases, the black hole stops exchanging radiation with its surroundings and becomes stable with a remnant mass. Then, we derived the equation of state function and discussed the pressure isotherms. Next, we investigated the shadow casts through Carter’s approach, We obtained the null geodesic equations to describe the photon motion. Then, we found the effective potential and showed the impacts of dark matter and quintessence matter fields on it. We demonstrated the shadows in celestial coordinates. We found that dark matter significantly lessens the shadow radius, while quintessence matter does not. Finally, we studied the quasinormal modes with the WKB and Mashhoon approximations. We tabulated quasinormal frequencies and showed the impact of dark matter and quintessence matter fields.

### v2

@@ 3 Black Hole’s Thermodynamics
Fig. 3 shows that the Hawking temperature decreases in all cases as the event horizon increases similar to the ordinary case. However, only for $\omega_{q}=-0.35$, the Hawking temperature tends to zero at infinite horizon radius as in the standard case. For the other three cases, the Hawking temperature $T_{H}$ decreases as the event horizon radius $r_{H}$ increases, eventually approaching zero for very large black holes. This implies that larger black holes become colder and emit less radiation. As $\alpha$ increases, the curves shift upward and rightward, indicating that the presence of PFDM increases the temperature for black holes of the same size. This effect is more pronounced for smaller black holes, as evidenced by the greater divergence of curves at smaller radii. However, as $r_{H}$ grows, the temperature converges to zero for all values of $\alpha$, suggesting that the effect of PFDM diminishes for more massive black holes. The behavior of the temperature approaching zero hints at a potential final state where black holes reach a stable, cold configuration, with no further Hawking radiation being emitted. This could suggest the formation of black hole remnants. For the larger values of the event horizon radius the temperature $T_{H}$ becomes negative. Recent experiments [202], have proven the existence of particles with negative temperatures in their motion. In [203], the author studied the possibility of such negative temperatures existing within black holes and proposed that this could generate an outward pressure counteracting the gravitational pull that draws matter inward.
@@
As we emphasized before, we see that the $\omega_{q}=-0.35$ scenario differs from the other cases. Fig. 5 shows that as $r_{H}$ increases, the heat capacity becomes more negative, suggesting larger black holes have a greater tendency towards instability [204]. Additionally, smaller values of the PFDM parameter $\alpha$ correspond to more negative heat capacities, implying that lesser dark matter presence leads to more pronounced black hole instability. In other cases, we observe that the heat capacity is first negative and then positive. This indicates a first-order phase transition at the horizon where specific heat equals zero. In other words, the vanishing value of the specific heat states that the black hole ceases to exchange radiation with its surroundings. Therefore, at that horizon black hole becomes stable and a remnant mass occurs. To clarify these discussions, we calculate the event horizon radii where specific heat disappears and present their values in Table 3.
@@ 4 Black Hole’s Shadow
in terms of $r_{ph}$. After we substitute the metric function, Eq. (18), we arrive at a subsequent equation
@@ 5.2 Mashhoon approximation
In [217], Mashhoon proposed to use Pöschl-Teller (PT) potential for calculating the quasinormal modes
@@
Here, $r_{0}^{\ast}$ is the position of the maximum of $\mathcal{V}\left(r^{\ast}\right)$ where $\mathcal{V}\left(r_{0}^{\ast}\right)=V_{0}$. The quasinormal modes of the PT potential can be computed analytically
@@ 6 Conclusions
In this paper, we handled the Schwarzschild black hole surrounded by PFDM in the presence of the quintessence matter field, and we explored its thermodynamics, shadows, and quasinormal modes. To this end, we initially derived the metric function of the black hole and discussed the metric function’s characteristics in four different quintessence field scenarios. We found that the quintessence matter field significantly alters the behavior of the metric function unless the quintessence state parameter is large enough. Then, we examined the mass function. We observed that dark matter sets a lower limit on the event horizon, but also expands its upper bound value. We noted that achieves a maximum mass value at a larger event horizon with a greater dark matter parameter. Next, we examined the black hole’s thermodynamics. At first, We derived the modified Hawking temperature. We found that it tends to zero unless the quintessence state parameter is not large enough. We showed the event horizon entropy is not affected by dark matter and quintessence matter field. Then. we investigated the heat capacity function. We showed that except very large quintessence state parameter case, the black hole has a first-order phase transition. In these cases, the black hole stops exchanging radiation with its surroundings and becomes stable with a remnant mass. Then, we derived the equation of state function and discussed the pressure isotherms. Next, we investigated the shadow casts through Carter’s approach, We obtained the null geodesic equations to describe the photon motion. Then, we found the effective potential and showed the impacts of dark matter and quintessence matter fields on it. We demonstrated the shadows in celestial coordinates. We found that dark matter significantly lessens the shadow radius, while quintessence matter does not. Finally, we studied the quasinormal modes with the WKB and Mashhoon approximations. We tabulated quasinormal frequencies and showed the impact of dark matter and quintessence matter fields.